# Progenitor diversity in the accreted stellar halos of Milky Way-like galaxies

### v1

@@ Abstract
Ongoing large stellar spectroscopic surveys of the Milky Way seek to reconstruct the major events in the assembly history of the Galaxy. Chemical and kinematic observations can be used to separate the contributions of different progenitor galaxies to the present-day stellar halo. Here we compute the number of progenitors that contribute to the accreted stellar halos of simulated Milky Way-like galaxies as a function of radius (the radial diversity) in three suites of models: Bullock & Johnston, Aquarius and Auriga. We show that there are significant differences between the predictions of these three models, beyond the halo-to-halo scatter expected in $\Lambda$CDM. Predictions of diversity from numerical simulations are sensitive to model-dependent assumptions regarding the efficiency of star formation in dwarf galaxies. We compare, at face value, to current constraints on the radial diversity of the Milky Way’s accreted halo. These constraints imply that the halo of our Galaxy is dominated by $\sim 2$ progenitors in the range $8-45$ kpc, in contrast to averages of $7$ progenitors in the Bullock & Johnston models, $3.5$ in Aquarius and $4.2$ in Auriga over the same region. We additionally find that the models with radial diversity most similar to that of the Milky Way are predominantly those with ongoing merger events. The Milky Way therefore appears unusual in having an accreted stellar halo dominated by a small number of progenitors accreted at very early times.
@@ 1 Introduction
The tidal disruption of dwarf galaxies is thought to be the most important process responsible for building up diffuse, metal-poor stellar halos around galaxies like the Milky Way (Searle & Zinn, 1978; Johnston et al., 1996). Clear evidence for this process is provided by the many stellar halo substructures that have been discovered in the Milky Way itself (e.g. Helmi, 2020; Bonaca & Price-Whelan, 2024), around M31 (e.g. Ibata et al., 2014a), and around other nearby galaxies (e.g. Martínez-Delgado et al., 2010). Some of these are associated with ongoing disruption of a surviving remnant, most notably the Sagittarius stream (Ibata et al., 1995; Newberg et al., 2002; Majewski et al., 2003). The Gaia satellite has enabled rapid progress in understanding the number, nature and origin of the progenitor galaxies that have contributed to the Milky Way’s stellar halo (Helmi, 2020; Deason & Belokurov, 2024). In particular, many prominent structures have been discovered in phase space that were not readily apparent in configuration space, including the Gaia Sausage-Enceladus feature (GS/E, Haywood et al., 2018; Helmi et al., 2018; Belokurov et al., 2018a). These heavily phase-mixed debris components likely correspond to the most massive and earliest accretion events (Belokurov et al., 2018b), which are expected to be relatively centrally concentrated in the Galactic potential (e.g. Johnston et al., 2008; Amorisco, 2017; Orkney et al., 2022). They can be identified as overdensities of stars with similar chemical abundances and orbital properties such as energy, angular momentum or actions (e.g. Helmi & de Zeeuw, 2000; McMillan & Binney, 2008; Gómez et al., 2010; Nissen & Schuster, 2010; Myeong et al., 2018a). Massive progenitors can also be identified by similar chemodynamical analyses of globular clusters (Lynden-Bell & Lynden-Bell, 1995; Myeong et al., 2018b; Kruijssen et al., 2020). Chemical abundance and radial velocity measurements from ongoing deep and wide spectroscopic surveys will likely lead to more discoveries and better characterization of features already known (Conroy et al., 2019; de Jong et al., 2019; Cooper et al., 2023; Jin et al., 2023). Nevertheless, current data are already sufficient to make quantitative estimates of the fraction of mass contributed to the stellar halo by its few most massive progenitors (e.g. Naidu et al., 2020, 2021), albeit still with substantial uncertainty around how those progenitors are identified (e.g. Carrillo et al., 2024) and how local observations are extrapolated (e.g. Sharpe et al., 2024). These successes and challenges motivate direct comparison between the available data and the predictions of simulations.
@@
In general, cosmological simulations of stellar halos associated with Milky Way-like galaxies appear to be in reasonable qualitative agreement with the observed mass, density, chemical composition and kinematics of the real Milky Way stellar halo (Bullock et al., 2000; Bullock & Johnston, 2005; Cooper et al., 2010; Deason et al., 2016; Monachesi et al., 2016b, 2019; Font et al., 2020). Agreement in this context means that the (often highly uncertain) observed properties lie within the (considerable) system-to-system and model-to-model scatter predicted by the simulations for isolated systems of present-day virial mass $0.5\lesssim M_{\mathrm{vir}}\lesssim 2\,\times 10^{12}\,\mathrm{M_{\odot}}$. Different simulations also agree on the most significant factors that determine the bulk properties of stellar halos, including the efficiency of galaxy formation in low-mass progenitor at high redshift, and the effectiveness of tidal disruption (dependent on the structure both of the progenitors and the central galaxy). Simulations may also produce a halo-like component through in situ star formation, unrelated (or only indirectly related) to galactic accretion (Abadi et al., 2003; Tissera et al., 2013; Cooper et al., 2015; Monachesi et al., 2019; Font et al., 2020). The presence of an ‘in situ’ stellar halo is perhaps the most significant source of potential ‘confusion’ in attempts to reconstruct the assembly history of the Milky Way (e.g. Rey & Starkenburg, 2022).
@@
Prior work on the composition of stellar halos has established that a small number of the most massive disrupted progenitors most likely account for the bulk of the total stellar halo mass. We discuss this more quantitatively in section 3.1. In a very broad sense, this a readily predictable consequence of the steep decline of star formation efficiency towards lower virial mass, combined with the characteristic accretion histories of Milky Way analogues (Bullock et al., 2000; Purcell et al., 2007; Bullock & Johnston, 2005; Cooper et al., 2010, 2013; Deason et al., 2013, 2016; Monachesi et al., 2019; D’Souza & Bell, 2018). Generally, the most-massive progenitors correspond to early mergers between the Milky Way analogue and systems with similar mass at the time of the merger (e.g. Cooper et al., 2010; Horta et al., 2024). The accretion of similarly massive systems at later times is relatively rare. There is some evidence suggesting that the Milky Way may be an example of this scenario, in which a small number of early-accreted massive progenitors dominate the total stellar halo mass (e.g. Deason et al., 2013), although the bulk of the mass from such progenitors is expected to be deposited close to the center of the halo, largely out of reach of current surveys of the Milky Way. This picture has received substantial support from observations of Milky Way analogues. In particular, where individual halo stars can be resolved, the tight galactic mass-metallicity relation enables the total stellar mass of the most massive progenitor to be estimated (to an accuracy of $\lesssim 1$ dex) from the peak of the metallicity distribution function (e.g. Harmsen et al., 2017; Bell et al., 2017; Jang et al., 2020; Gozman et al., 2023). This can be combined with an estimate of the total accreted mass (e.g. from fits to the outer regions of the surface brightness profile) to determine the extent to which the most massive progenitor dominates; this approach has been validated in cosmological simulations (e.g D’Souza & Bell, 2018; Monachesi et al., 2019). Observational results from this approach are generally in line with the expectation above, that $\lesssim 10$ progenitors account for the bulk of the accreted mass, with a large fraction contributed by the most massive progenitor. Using the Auriga simulations, Monachesi et al. (2019) find that the radial gradient of metallicity can, in some cases, further constrain the number of less massive progenitors that make significant contributions to the halo.
@@
The generic expectation described above – that a small number of progenitors will dominate the accreted stellar halo at the virial mass scale of the Milky Way – relates to the entirety of the accreted mass. In practice, particularly in the Milky Way, observations can presently only sample a small fraction of the total volume of the halo, which may lead to different conclusions about the nature and number of the most significant progenitors (e.g. Sharpe et al., 2024). The relative significance of different progenitors may vary with radius (Cooper et al., 2010; Deason et al., 2013, 2014, 2016; Amorisco, 2017; Fattahi et al., 2020). It is therefore relevant to ask if simulations predict that the region of the accreted stellar halo probed by current surveys of the Milky Way (i.e. within heliocentric distances of $\lesssim 20$ kpc) is dominated by one object or many, and whether the inferred masses and accretion times of the significant progenitors inferred from the Milky Way data are consistent with expectations for a ‘typical’ $10^{12}\,\mathrm{M_{\odot}}$ Milky Way host.
@@
Our aim in this paper is to explore the radial dependence of the ‘diversity’ of progenitor contributions in Milky Way analogue stellar halos using three commonly cited suites of Milky Way stellar halo simulations: the semi-analytic/hybrid $N$-body model of Bullock & Johnston (2005), the semi-analytic particle-tagging Aquarius models of Cooper et al. (2010) and the hydrodynamical Auriga models of Grand et al. (2017). We compare these predictions, at face value, with the inferences made from the current data by Naidu et al. (2020), which are broadly representative of recent interpretations of Milky Way observations. The simulations differ in their numerical approach and assumptions regarding the physics of galaxy formation. Our goal is to summarise the systematic differences between their predictions, in the specific contexts in which they are most readily compared to current stellar halo observations. To support our comparison, we also make available the underlying simulation data of Cooper et al. (2010).
@@
In detail, this kind of comparison is best conducted with mock observables (e.g. Lowing et al., 2015; Grand et al., 2018; Sanderson et al., 2018; Kizhuprakkat et al., 2024), rather than directly from the ‘raw’ simulation data, in order to account for observational biases and systematic uncertainties, for example in total densities estimated from small samples of tracers. Each of the three simulations we examine has been used to make publicly available mock star catalogs, which could be used for this purpose. Here, however, we use the raw simulations because our aim is only to establish a ‘baseline’, for the range of underlying predictions from the models, independent of the parameters and selection function of any particular survey. We hope this will be useful to underpin more detailed future comparisons between specific surveys and mock catalogues.
@@
The paper is structured as follows. In Section 2, we briefly summarize the assumptions, methods, initial conditions, and relevant results from each of the three simulation suites (2.1 to 2.3). We compare their predictions for accreted stellar halo masses and density profiles in section 2.4. In Section 3 we present their radial diversity profiles. We summarize our findings and conclude in Section 4.
@@ 2 Simulations
In this section we briefly describe the three sets of simulations and compare their predictions for accreted stellar halos. Table 1 lists the cosmological parameters assumed by each model. Each model assumes a slightly different cosmology, but we do not attempt to homogenize them in our comparison. In particular, where the simulations provide quantities scaled to $H_{0}=100\,h\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, we convert these to ‘observed’ quantities using the value of $h$ from the cosmological parameter set assumed by the simulation (i.e. we use different values of $h$ for each simulation, as in Table 1).
@@ 2.1 Bullock & Johnston simulations
Notwithstanding these limitations, the BJ05 models (supplemented by Robertson et al., 2005; Font et al., 2006) have long served as the most widely-used and influential set of numerical predictions for the phenomenology of galactic stellar halos, and for interpreting observations of the Milky Way and M31 (e.g. Sharma et al., 2011; Bell et al., 2008; Johnston et al., 2008; Ibata et al., 2014b; Sharpe et al., 2024). Among their many results, the BJ05 models were the first to clearly demonstrate the principle discussed in the introduction, that the bulk of the mass in the stellar halo of a typical Milky Way analogue can be (and often is) deposited by a relatively small number of satellite progenitors with total mass $M_{\mathrm{vir}}\sim 10^{10}\,\mathrm{M_{\odot}}$. BJ05 quantify this by the fraction of accreted stellar halo mass associated with the 15 most massive progenitors, which they find to be $\gtrsim 80~{}\%$ in all 11 halos (their Table 1). Robertson et al. (2005) and Font et al. (2006) established that the chemical abundances observed for Milky Way halo stars further support growth from a small number of massive progenitors, and moreover, that those progenitors must have been accreted relatively early, $>9$ Gyr ago.
@@ 2.2 Aquarius
Aquarius is a suite of six high-resolution cosmological N-body ‘zoom’ re-simulations of Milky Way-like dark matter halos (Springel, 2010). Cooper et al. (2010, hereafter C10) processed merger trees for the ‘level 2’ Aquarius simulations with the semi-analytic galaxy formation model galform (Cole et al., 2000; Bower et al., 2006; Font et al., 2011). The star formation histories predicted by galform were used as the basis for a particle tagging technique, which associates single stellar populations with subsets of tightly-bound dark matter particles in a ‘continuous’ fashion along each branch of the merger tree. Using the tagged particles, the evolving 3-dimensional spatial distribution and kinematics of each stellar population can be tracked through the $N$-body simulation in post-processing, under the assumption that the baryonic mass does not significantly perturb the potential as represented by the underlying $N$-body model. This procedure is described in detail by Cooper et al. (2017), along with a study of its limitations and correspondence to the results of hydrodynamical methods. The properties of each Aquarius stellar halo from C10 are summarised in Appendix A. With this paper, we make available the particle data from the C10 models (see the data availability statement).
In C10, in contrast to BJ05, the evolution in the mass and structure of the main halo and its satellites are calculated ab initio; in this $N$-body sense the Aquarius simulations are ‘fully cosmological’ and dynamically self-consistent. The central potential is non-spherical, and can grow violently as well as through smooth accretion. Aquarius is therefore a more realistic model than BJ05 at early times, prior to the last merger that strongly perturbs the central potential. Moreover, the Aquarius sample is not restricted to particular assembly histories, and hence, unlike BJ05, includes halos that grow significantly through the accretion of relatively massive progenitors at $z<1$. Nevertheless, the particle-tagging scheme involves significant approximations. In particular, it does not account for the growth of a massive baryonic disk in the main halo, which may have a substantial effect on the disruption of satellites, both those accreted at early times and massive satellites accreted relatively late (Cooper et al., 2017).
@@
C10 further separated the accreted component into two regions: “bulge” (galactic radius $R<3\,\mathrm{kpc}$) and ‘stellar halo’ ($3<R<200\,\mathrm{kpc}$). Here we study the properties of accreted stars as a function of radius, so we treat the $R<3\,\mathrm{kpc}$ region in the same way as the rest of the accreted component.

### v2

# Progenitor diversity in the accreted stellar halos of Milky Way–like galaxies

@@ Abstract
Ongoing large stellar spectroscopic surveys of the Milky Way seek to reconstruct the major events in the assembly history of the Galaxy. Chemical and kinematic observations can be used to separate the contributions of different progenitor galaxies to the present-day stellar halo. Here, we compute the number of progenitors that contribute to the accreted stellar halos of simulated Milky Way–like galaxies as a function of radius (the radial diversity) in three suites of models: Bullock & Johnston, Aquarius, and Auriga. We show that there are significant differences between the predictions of these three models, beyond the halo-to-halo scatter expected in $\Lambda$CDM. Predictions of diversity from numerical simulations are sensitive to model-dependent assumptions regarding the efficiency of star formation in dwarf galaxies. We compare, at face value, to current constraints on the radial diversity of the Milky Way’s accreted halo. These constraints imply that the halo of our Galaxy is dominated by $\sim 2$ progenitors in the range 8–45 kpc, in contrast to averages of $7$ progenitors in the Bullock & Johnston models, $3.5$ in Aquarius, and $4.2$ in Auriga over the same region. We additionally find that the models with radial diversity most similar to that of the Milky Way are predominantly those with ongoing merger events. The Milky Way therefore appears unusual in having an accreted stellar halo dominated by a small number of progenitors accreted at very early times.
@@ 1 Introduction
The tidal disruption of dwarf galaxies is thought to be the most important process responsible for building up diffuse, metal-poor stellar halos around galaxies like the Milky Way (Searle & Zinn, 1978; Johnston et al., 1996). Clear evidence for this process is provided by the many stellar halo substructures that have been discovered in the Milky Way itself (e.g. Helmi, 2020; Bonaca & Price-Whelan, 2025), around M31 (e.g. Ibata et al., 2014), and around other nearby galaxies (e.g. Martínez-Delgado et al., 2010). Some of these are associated with ongoing disruption of a surviving remnant, most notably the Sagittarius stream (Ibata et al., 1995; Newberg et al., 2002; Majewski et al., 2003). The Gaia satellite has enabled rapid progress in understanding the number, nature, and origin of the progenitor galaxies that have contributed to the Milky Way’s stellar halo (Helmi, 2020; Deason & Belokurov, 2024). In particular, many prominent structures have been discovered in phase space that were not readily apparent in configuration space, including the Gaia Sausage–Enceladus feature (GS/E, Haywood et al., 2018; Helmi et al., 2018; Belokurov et al., 2018). These heavily phase-mixed debris components likely correspond to the most massive and earliest accretion events (Belokurov et al., 2018), which are expected to be relatively centrally concentrated in the Galactic potential (e.g. Johnston et al., 2008; Amorisco, 2017a; Orkney et al., 2022). They can be identified as overdensities of stars with similar chemical abundances and orbital properties such as energy, angular momentum, or actions (e.g. Helmi & de Zeeuw, 2000; McMillan & Binney, 2008; Gómez et al., 2010; Nissen & Schuster, 2010; Myeong et al., 2018a). Massive progenitors can also be identified by similar chemodynamical analyses of globular clusters (Lynden-Bell & Lynden-Bell, 1995; Myeong et al., 2018b; Kruijssen et al., 2020). Chemical abundance and radial velocity measurements from ongoing deep and wide spectroscopic surveys will likely lead to more discoveries and better characterization of features already known (Conroy et al., 2019; de Jong et al., 2019; Cooper et al., 2023; Jin et al., 2023). Nevertheless, current data are already sufficient to make quantitative estimates of the fraction of mass contributed to the stellar halo by its few most massive progenitors (e.g. Naidu et al., 2020, 2021), albeit still with substantial uncertainty around how those progenitors are identified (e.g. Lane et al., 2023; Carrillo et al., 2024) and how local observations are extrapolated (e.g. Sharpe et al., 2024). These successes and challenges motivate direct comparison between the available data and the predictions of simulations.
@@
In general, cosmological simulations of stellar halos associated with Milky Way–like galaxies appear to be in reasonable qualitative agreement with the observed mass, density, chemical composition, and kinematics of the real Milky Way stellar halo (Bullock et al., 2001; Bullock & Johnston, 2005; Cooper et al., 2010; Deason et al., 2016; Monachesi et al., 2016b, 2019; Font et al., 2020). Agreement in this context means that the (often highly uncertain) observed properties lie within the (considerable) system-to-system and model-to-model scatter predicted by the simulations for isolated systems of present-day virial mass $0.5\lesssim M_{\mathrm{vir}}\lesssim 2\,\times 10^{12}\,\mathrm{M_{\odot}}$. Different simulations also agree on the most significant factors that determine the bulk properties of stellar halos, including the efficiency of galaxy formation in low-mass progenitor at high redshift, as well as the effectiveness of tidal disruption (dependent on the structure both of the progenitors and the central galaxy). Simulations may also produce a halo-like component through in situ star formation, unrelated (or only indirectly related) to galactic accretion (Abadi et al., 2003; Tissera et al., 2013; Cooper et al., 2015; Monachesi et al., 2019; Font et al., 2020). The presence of an ‘in situ’ stellar halo is perhaps the most significant source of potential ‘confusion’ in attempts to reconstruct the assembly history of the Milky Way (e.g. Rey & Starkenburg, 2022). Chemical and kinematic evidence points to an in situ contribution to the Milky Way halo, at least in the inner galaxy, in particular from stars formed in the proto-Milky Way and later scattered onto eccentric orbits (e.g. Nissen & Schuster, 2010; Gallart et al., 2019; Belokurov et al., 2020).
Prior work on the composition of stellar halos has established that a small number of the most massive disrupted progenitors most likely account for the bulk of the total stellar halo mass. We discuss this more quantitatively in section 3.1. In a very broad sense, this a readily predictable consequence of the steep decline of star formation efficiency toward lower virial mass, combined with the characteristic accretion histories of Milky Way analogs (Bullock et al., 2001; Purcell et al., 2007; Bullock & Johnston, 2005; Cooper et al., 2010, 2013; Deason et al., 2013, 2016; Monachesi et al., 2019; D’Souza & Bell, 2018).
A range of scenarios for the assembly of the accreted stellar halo arise from the range of CDM halo mass growth histories at a fixed present-day virial mass (e.g. Amorisco, 2017b). The picture is more complicated than might be expected based on predictions for satellite merger rates alone, because satellites must be heavily disrupted in order to become significant stellar halo progenitors. On average, satellites with the highest total stellar mass arrive late, and are less likely to be disrupted by $z=0$. For a Milky Way analog, the dominant stellar halo progenitors are therefore typically less massive than the most massive surviving satellite, and early-arriving massive progenitors can make relatively significant contributions. This suggests a complex but potentially informative relationship between the assembly history of a galaxy and the mass, diversity, and structure of its accreted stellar halo. For example, Amorisco (2017b) show that, on average, the total mass of the accreted stellar halo is maximized when a relatively larger number of progenitors of moderate mass ($M_{\star}\sim 10^{8}\,\mathrm{M_{\odot}}$) arrive at intermediate redshift ($z\sim 1.5$). This results in a high stellar mass per progenitor and allows enough time to incorporate those satellites into the stellar halo. At fixed present-day virial mass, such an assembly history requires the system to lag behind the average growth rate at earlier and later times.
These expectations from simplified models are reflected in cosmological simulations, although the stochastic nature of massive accretion events can obscure trends in small samples. In general, simulated accreted halos of Milky Way analogs typically comprise a few substantial contributions from mergers at early times, along with a further one or two massive accretions of similar or greater stellar mass at late times (e.g. Bullock & Johnston, 2005; Cooper et al., 2010; D’Souza & Bell, 2021; Horta et al., 2024).
There is some evidence suggesting that the Milky Way may be an example of a system assembled rapidly and early, in which a small number of early-arriving moderately massive progenitors dominate the total accreted stellar halo mass (e.g. Deason et al., 2013). However, the bulk of the mass from such progenitors is expected to be deposited close to the center of the halo (e.g. Amorisco, 2017a), largely out of reach of current surveys of the Milky Way.
@@
Observations of Milky Way analogs are also in line with the expectation that $\lesssim 10$ progenitors account for the bulk of the accreted mass, with a large fraction contributed by the most massive progenitor. In particular, where individual halo stars can be resolved, the tight galactic mass-metallicity relation enables the total stellar mass of the most massive progenitor to be estimated (to an accuracy of $\lesssim 1$ dex) from the peak of the metallicity distribution function (e.g. Harmsen et al., 2017; Bell et al., 2017; Jang et al., 2020; Gozman et al., 2023). This can be combined with an estimate of the total accreted mass (e.g., from fits to the outer regions of the surface brightness profile) to determine the extent to which the most massive progenitor dominates; this approach has been validated in cosmological simulations (e.g., D’Souza & Bell, 2018; Monachesi et al., 2019). Using the Auriga simulations, Monachesi et al. (2019) find that the radial gradient of metallicity can, in some cases, further constrain the number of less massive progenitors that make significant contributions to the halo.
@@
The generic expectation described above – that a small number of progenitors will dominate the accreted stellar halo at the virial mass scale of the Milky Way – relates to the entirety of the accreted mass. In practice, in particular in the Milky Way, observations can presently only sample a small fraction of the total volume of the halo, which may lead to different conclusions about the nature and number of the most significant progenitors (e.g. Sharpe et al., 2024). The relative significance of different progenitors may vary with radius (Cooper et al., 2010; Deason et al., 2013, 2014, 2016; Amorisco, 2017a; Fattahi et al., 2020). It is therefore relevant to ask if simulations predict that the region of the accreted stellar halo probed by current surveys of the Milky Way (i.e., within heliocentric distances of $\lesssim 20$ kpc) is dominated by one object or many, and whether the inferred masses and accretion times of the significant progenitors inferred from the Milky Way data are consistent with expectations for a ‘typical’ $10^{12}\,\mathrm{M_{\odot}}$ Milky Way host.
@@
Our aim in this paper is to explore the radial dependence of the ‘diversity’ of progenitor contributions in Milky Way analog stellar halos using three commonly cited suites of Milky Way stellar halo simulations: the semi-analytic/hybrid $N$-body model of Bullock & Johnston (2005), the semi-analytic particle-tagging Aquarius models of Cooper et al. (2010) and the hydrodynamical Auriga models of Grand et al. (2017). We compare these predictions, at face value, with the inferences made from the current data by Naidu et al. (2020), which are broadly representative of recent interpretations of Milky Way observations. The simulations differ in their numerical approach and assumptions regarding the physics of galaxy formation. Our goal is to summarize the systematic differences between their predictions, in the specific contexts in which they are most readily compared to current stellar halo observations. To support our comparison, we also make available the underlying simulation data of Cooper et al. (2010).
@@
In detail, this kind of comparison is best conducted with mock observables (e.g. Lowing et al., 2015; Grand et al., 2018; Sanderson et al., 2020; Kizhuprakkat et al., 2024), rather than directly from the ‘raw’ simulation data, in order to account for observational biases and systematic uncertainties, for example in total densities estimated from small samples of tracers. Each of the three simulations we examine has been used to make publicly available mock star catalogs, which could be used for this purpose. Here, however, we use the raw simulations because our aim is only to establish a ‘baseline’, for the range of underlying predictions from the models, independent of the parameters and selection function of any particular survey. We hope this will be useful to underpin more detailed future comparisons between specific surveys and mock catalogs.
@@
The paper is structured as follows. In Section 2, we briefly summarize the assumptions, methods, initial conditions, and relevant results from each of the three simulation suites (Sections 2.1 to 2.3). We compare their predictions for accreted stellar halo masses and density profiles in section 2.4. In Section 3 we present their radial diversity profiles. We summarize our findings and conclude in Section 4.
@@ 2 Simulations
In this section, we briefly describe the three sets of simulations and compare their predictions for accreted stellar halos. Table 1 lists the cosmological parameters assumed by each model. Each model assumes a slightly different cosmology, but we do not attempt to homogenize them in our comparison. In particular, where the simulations provide quantities scaled to $H_{0}=100\,h\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, we convert these to ‘observed’ quantities using the value of $h$ from the cosmological parameter set assumed by the simulation (i.e. we use different values of $h$ for each simulation, as in Table 1).
@@ 2.1 Bullock & Johnston simulations
Notwithstanding these limitations, the BJ05 models (supplemented by Robertson et al., 2005; Font et al., 2006) have long served as the most widely used and influential set of numerical predictions for the phenomenology of galactic stellar halos, and for interpreting observations of the Milky Way and M31 (e.g., Bell et al., 2008; Johnston et al., 2008; Sharma et al., 2011; Ibata et al., 2014; Sharpe et al., 2024). Among their many results, the BJ05 models were the first to clearly demonstrate the principle discussed in the introduction, i.e., that the bulk of the mass in the stellar halo of a typical Milky Way analog can be (and often is) deposited by a relatively small number of satellite progenitors with total mass $M_{\mathrm{vir}}\sim 10^{10}\,\mathrm{M_{\odot}}$. BJ05 quantify this by the fraction of accreted stellar halo mass associated with the 15 most massive progenitors, which they find to be $\gtrsim 80~{}\%$ in all 11 halos (their Table 1). Robertson et al. (2005) and Font et al. (2006) established that the chemical abundances observed for Milky Way halo stars further support growth from a small number of massive progenitors, and moreover, that those progenitors must have been accreted relatively early, $>9$ Gyr ago.
@@ 2.2 Aquarius
Aquarius is a suite of six high-resolution cosmological $N$-body ‘zoom’ resimulations of Milky Way–like dark matter halos (Springel, 2010). Cooper et al. (2010, hereafter C10) processed merger trees for the ‘level 2’ Aquarius simulations with the semi-analytic galaxy formation model galform (Cole et al., 2000; Bower et al., 2006; Font et al., 2011). The star formation histories predicted by galform were used as the basis for a particle-tagging technique, which associates single stellar populations with subsets of tightly bound dark matter particles in a ‘continuous’ fashion along each branch of the merger tree. Using the tagged particles, the evolving three-dimensional spatial distribution and kinematics of each stellar population can be tracked through the $N$-body simulation in post-processing, under the assumption that the baryonic mass does not significantly perturb the potential as represented by the underlying $N$-body model. This procedure is described in detail by Cooper et al. (2017), along with a study of its limitations and correspondence to the results of hydrodynamical methods. The properties of each Aquarius stellar halo from C10 are summarized in the Appendix A. With this paper, we make available the particle data from the C10 models (see the data availability statement).
In C10, in contrast to BJ05, the evolution in the mass and structure of the main halo and its satellites are calculated ab initio; in this $N$-body sense, the Aquarius simulations are ‘fully cosmological’ and dynamically self-consistent. The central potential is nonspherical, and it can grow violently as well as through smooth accretion. Aquarius is therefore a more realistic model than BJ05 at early times, prior to the last merger that strongly perturbs the central potential. Moreover, the Aquarius sample is not restricted to particular assembly histories, and hence, unlike BJ05, it includes halos that grow significantly through the accretion of relatively massive progenitors at $z<1$. Nevertheless, the particle-tagging scheme involves significant approximations. In particular, it does not account for the growth of a massive baryonic disk in the main halo, which may have a substantial effect on the disruption of satellites, both those accreted at early times and massive satellites accreted relatively late (Cooper et al., 2017).
@@
C10 further separated the accreted component into two regions: ‘bulge’ (galactic radius $R<3\,\mathrm{kpc}$) and ‘stellar halo’ ($3<R<200\,\mathrm{kpc}$). Here, we study the properties of accreted stars as a function of radius, so we treat the $R<3\,\mathrm{kpc}$ region in the same way as the rest of the accreted component.